OpenStudy (anonymous):

solve using the quadratic equation by factoring and applying the zero product property. 5x^2=-8x

OpenStudy (anonymous):

i got 5x^2-8x=0

OpenStudy (anonymous):

x(5x-8)

OpenStudy (anonymous):

x(5x-8) = 0 so either x = 0 OR 5x-8 = 0